# This Candidate is [MASK]

### v1

@@ 2.2 Prompt-based Learning
In recent years, there has been a paradigm shift in natural language processing (NLP), inspired by the development and fast popularization of large pre-trained transformer-based language models (PLMs). In a nutshell, the paradigm shift involved moving from a pre-train then fine-tune strategy to the prompt-based learning approach. [^14] For the sake of brevity, I limit the description of the paradigm shift and of prompt-based learning to the bare minimum. See Sun et al. (2022), Min et al. (2023) and Liu et al. (2023) for comprehensive reviews. Bu et al. (2024) focus on sentiment analysis via prompt learning.
The idea of pre-training is to first train one large model on a “fundamental” pre-training task and only later adapt (“fine-tune”) it to a multiplicity of specialized tasks in a second step. In the context of NLP, the pre-training task has always been some kind of language modeling, which involves either unidirectionally predicting the next word in a sequence of text, or a fill-in-the-blank style task in which the model must predict a masked word in a sentence. [^15] The ULMFiT model (Howard and Ruder, 2018) is widely credited with popularizing pre-training and fine-tuning for NLP tasks. That model pre-trained by predicting the next token in a sequence. The great advantage of pre-training is that the model can exploit the massive quantities of text available from the web and other sources, largely obviating the extremely expensive labeling required in classical supervised learning.
The pre-training technique has been instrumental in the development of large language models following the development of the transformer architecture (Vaswani et al., 2017). The transformer’s attention mechanism allows the processing of text sequences without the computationally expensive recurrent processing required in LSTMs and the other recurrent neural network layers that were the previous state of the art. [^16] See Zhao et al. (2023) for a detailed review of the recent advances of LLMs. Importantly, soon after the first transformers were rolled out, it became apparent to researchers that the models had capabilities that far exceeded simple language generation. The newly discovered “emergent properties” of LLMs range from text classification to translation tasks (see Brown et al., 2020).
@@
Instead of fine-tuning a PLM to downstream tasks, these tasks can be reformulated to look more like those solved during the model’s pre-training. The key insight is that a textual prompt can be devised to elicit the desired prediction. For example, in order to recognize the emotion of the tweet “I missed the bus today!”, we may append the prompt “I felt so [MASK]” and ask the LLM to fill the blank with an emotion-bearing word. [^17] The same prompting strategy can be applied to multiple tasks. To elicit a translation, one might use the prompt: “English: I missed the bus today! French: …” and use a generative LLM to generate the translation.
@@
The great advantage of reformulating the problem this way is that, once the prompting strategy has been set, one may proceed to making predictions without the need of any further training. Specifically, prompting spares us from the often unsurmountable task of collecting a large labeled data set. Moreover, in many cases —including assigning sentiment to confidential reference letters— the difficulty inherent in the labeling process itself may become an obstacle.
To give a formal statement of these points, consider first a supervised learning NLP task that takes a text $\boldsymbol{x}$ and predicts output based on a model $\mathrm{P}\left(y\mid\boldsymbol{x};\boldsymbol{\theta}\right)$. For a text classification task, given an input text $\boldsymbol{x}$ the model predicts a label $y$ from a fixed label set $\mathcal{Y}$. For example, a sentiment analysis task might take an input such as “I loved this movie!” and predict a label $y=+$ out of the label set $\mathcal{Y}=\{+;\sim;-\}$.
In this terms, the challenge of supervised learning is that in order to find optimal weights $\boldsymbol{\theta}$, we need lots of supervised data $\left(\boldsymbol{x}_{i},y_{i}\right),i=1,\ldots,N$ where the $y_{i}$ are the labels (a.k.a. “ground truth”) for a number of training examples.
@@
Prompt-based learning takes advantage of a pre-trained model $\mathrm{P}\left(\boldsymbol{x};\boldsymbol{\gamma}\right)$ of text $\boldsymbol{x}$ itself. The derivation of predicted probabilities requires three steps:
@@
[…]
@@
The third and final step is simply to collect predicted probabilities for each label using the PLM. Let the function $f_{fill}(\boldsymbol{x}^{\prime},\boldsymbol{m})$ fill-in the location [MASK] in prompt $\boldsymbol{x}^{\prime}$ with the permissible answer $\boldsymbol{m}$. The predicted probabilities for each label are thus defined:
$\mathrm{P}(y\mid\boldsymbol{x})=$ $\sum_{\mathbf{m}\in\mathcal{M}}\mathrm{I}\left[\mathcal{V}(\boldsymbol{m})=y\right]\cdot\mathrm{P}\left(f_{fill}(f_{prompt}(\boldsymbol{x}),\boldsymbol{m});\boldsymbol{\gamma}\right),\quad y\in\mathcal{Y}$ (1)
where $\mathrm{I}[\cdot]$ is the indicator function. Intuitively, the prompt-based probabilities are a measure of how likely the PLM judges each of the filled prompts. In particular, the predicted probability of assigning positive sentiment to the input text is simply the sum of the probabilities — based on the LLM’s pre-training — assigned to each of the positive tokens in the verbalizer. This intuitive interpretation is one of the advantages of prompt learning. In addition, because sentiment scores are derived from the predicted probabilities, they have a natural unit of measurement.
@@
$\mathrm{polarity}(\boldsymbol{x})=$ $\mathrm{P}\left(y=+\mid\boldsymbol{x}\right)-\mathrm{P}\left(y=-\mid\boldsymbol{x}\right)$
2.3 Applying Prompt-learning to Reference Letters
In principle, any PLM capable of evaluating $\mathrm{P}\left(\boldsymbol{x};\boldsymbol{\gamma}\right)$ can be used for prompting. However, the model choice has important implications for the overall quality and precision of the sentiment scores, as well as for other somewhat more technical aspects.
@@ 2.3 Applying Prompt-learning to Reference Letters
    Generative models such as those in the GPT and Llama families require a prefix prompt. In other words, the [MASK] term in the template must be the last token. [^19] Because of this restriction, the [MASK] term is often omitted altogether. The reason for this is that generative models are pre-trained using causal language modeling, i.e. predicting the next word in text sequences. In contrast, bidirectional models like Bert, Roberta, or Big Bird are pre-trained via masked language modeling, so the [MASK] term can be anywhere in the template. [^20] These are called cloze prompts.
@@
    An important consideration is that the length of the input text (measured in tokens) must fit within the model’s context size, which varies from model to model.
In this paper, I use Big Bird (Zaheer et al., 2020), a bidirectional LLM pre-trained on Wikipedia, books, short stories, and news articles. Big Bird’s vocabulary contains over 50 thousand tokens, including almost every common-use English word. [^22] Big Bird’s tokenizer is based on the SentencePiece algorithm, so some tokens in the vocabulary are common word parts or punctuation.
Big Bird implements a special attention layer (named “sparse attention” by the Google team that developed it) with a context size of 4,096 tokens. The model can also be used in “full attention” mode, which reduces the context size to 1,024 tokens but increases performance. [^23] Because many documents in my data set exceed the 1,024 token limit, using the model in full attention mode requires applying a technique known as windowing. More details are provided below.
2.4 Sentiment Prompt and Verbalizer
To extract sentiment from reference letters, I use the following template/verbalizer combination:
$\mathcal{T}_{sentiment}:=$ “[X] In summary, this job market candidate is [MASK].”
$\mathcal{V}_{sentiment}:=$
    excellent, outstanding, exceptional, brilliant, great, superb, fantastic, terrific, amazing, phenomenal, awesome, stellar, wonderful, extraordinary, remarkable, impressive, incredible, fabulous, formidable, marvelous, perfect, fascinating.
    bad, terrible, disappointing, poor, awful, lousy, atrocious, pathetic, unsatisfactory, substandard, mediocre, common, inferior, average, ordinary, regular.
@@ 3 Hiring Data
The analysis in this paper is based on administrative data from my institution’s recruitment efforts during the 2013, and 2015–2021 job markets. [^24] I designate job market cycles by the year in which recruitment formally started and the job ads were posted. Unfortunately, data for the 2014 market was not archived. To protect confidentiality, I processed all data myself.
As already mentioned, I only consider job market candidates who are finishing their PhD programs and have not held a full-time academic position. In addition, only applications where the adviser’s letter of recommendation and at least one additional letter were submitted are kept. Overall, my data contains 1,998 letters from 603 job market candidates.
The pre-processing of the reference letters’ text was kept to a minimum. I removed headers, footers, page numbers, salutations, and some obvious mistakes from faulty conversions from pdf format. I also collected personal data from the submitted CVs, as well as personal websites and LinkedIn profiles. Specifically, I have data on the candidate’s sex, nationality, prior publications and time to PhD completion. I also collect the PhD granting institution’s location and ranking. [^25] To rank Economics departments, I use Tilburg University’s ranking, which is based on publications weighted by journal impact factor and covers the whole world (https://rankings.tilburguniversity.edu/).
Finally, I hand-collected the detailed job market outcome for each candidate in my data. The job market outcome corresponds to the job held by September of the following year.
@@ 3.1 Descriptive Statistics
Gender
Female 603 0.255 0 1
Citizenship Region
Developing Countries 603 0.093 0 1
Former Soviet Union 603 0.217 0 1
North America 603 0.036 0 1
Europe 603 0.347 0 1
Asia 603 0.307 0 1
PhD Region
USA 603 0.498 0 1
Europe 603 0.323 0 1
Other English Speaking 603 0.164 0 1
Other 603 0.015 0 1
Research Field
Finance 603 0.272 0 1
Applied 603 0.342 0 1
Macroeconomics 603 0.232 0 1
Theory 603 0.100 0 1
Econometrics 603 0.055 0 1
PhD Ranking
1–25 603 0.252 0 1
26–50 603 0.235 0 1
51–75 603 0.083 0 1
76–100 603 0.075 0 1
101–150 603 0.141 0 1
151–200 603 0.081 0 1
201+ 603 0.133 0 1
PhD Duration
Seven years or more 603 0.096 0 1
Publications
At Least One Top Field 603 0.053 0 1
Notes: “Other English Speaking” includes UK, Ireland Rep., Australia, and Singapore. The PhD ranking is from Tilburg. I considered publications in JIE, JET, JoE, JME, JPubE, JEH, JFE, JF, Rand, JEEA, REStat, EJ, IER, all the AEJs, and the top 5.
@@
Table 1 summarizes the main characteristics of the job market candidates in my data. There are some notable differences between my data and similar samples collected by Eberhardt et al. (2023) and Baltrunaite et al. (2024) from British and Italian institutions, respectively. Given that my institution is based in Russia, it is not surprising to see an over-representation of applicants from the former Soviet Union. [^26] Of these, 108 are Russian, 8 are Ukrainian, 5 are Armenian, 5 from the Baltic Countries, and 5 from other countries. There is also a relatively small fraction of North American citizens. About one in four applicants in my data are female, compared to around 30% in their samples. Another important difference is that the samples in their studies include both new PhDs and experienced academics. Because of this, the candidates in their samples have more publications.
@@
In other respects, the samples are comparable. In particular, the fraction of candidates getting their PhD from an American institutions (around 50%) and the fraction in top-ranked programs are very similar.
3.2 Reference Letter Polarity
Figure 1 shows the distribution of polarity scores and letter length for four different dictionary approaches: the Harvard General Inquirer, the Loughran-McDonald lexicon, Vader, and Senticnet. Several facts are apparent. First, there is no clear relationship between letter length and letter sentiment. A locally weighted regression line is practically flat in all cases.
Second, advisers’ letters are distinctly longer than other referees’ letters. On average, non-adviser letters are in the neighborhood of 900 words long. Advisers’ letters are almost 33% longer on average, and the difference is highly statistically significant. Finally, note how several of the lexicons find no sentiment-carrying words in a sizeable number of cases. These cases appear as a horizontal line of zero-polarity for HGI, Senticnet and Vader.
Figure 1: The Distribution of Polarity and Letter Length — Dictionary Approaches
Notes: Fitted line is a locally weighted regression. LM polarity re-scaled.
In figure 2, I present similar scatter plots using prompt-based sentiment extraction for different variants of the Big Bird Model. The rightmost plot depicts the result of using the default sparse attention, which is the main innovation of the Big Bird model (see Zaheer et al., 2020, for details). The great advantage of sparse attention is that the context window is expanded to 4,096 tokens, which is enough to fit all the reference letters in my data. Unfortunately, as evident in the figure, sparse attention introduces a strong negative dependence between letter length and sentiment polarity. The model is simply not able to cope with the amount of information present in the long letters.
Big Bird can also be used in full attention mode, which limits the context window to 1,024 tokens. In order to process longer letters, I applied a technique known as windowing. In a nutshell, windowing splits a long document into partially overlapping “windows”. After some experimentation, I came to the conclusion that setting the stride —i.e. overlap— to half the window size (512 tokens) worked well. [^27] I also fine tuned the algorithm so that text windows contained only full sentences.
Figure 2: The Distribution of Polarity and Letter Length — Big Bird Prompting
Notes: Fitted line is a locally weighted regression.
Table 2 shows the results of applying the windowing technique. Most letters can be covered with up to two windows but a few require up to seven windows. Interestingly, the results suggest that the highest sentiment content in letters of reference is located at the end of the document. Frequent reference letter readers will not be surprised. The only remaining issue is how to summarize the up to seven scores per letter. As shown in the middle plot of figure 2, simply averaging the scores predictably biases the sentiment polarity of longer letters downwards. Therefore, I use the maximum score over all windows —invariably, this corresponds to the last window— as my final measure of letter sentiment.
Table 2: Windowing Results
Total Window Average Polarity $N$
Windows w1 w2 w3 w4 w5 w6 w7
1 0.256 835
2 0.064 0.215 560
3 0.054 0.035 0.268 281
4 0.064 0.024 0.052 0.290 133
5 0.091 0.019 0.021 0.037 0.210 44
6 0.092 0.011 0.013 0.030 0.075 0.259 29
7 0.171 0.007 0.021 0.037 0.024 0.082 0.315 16
Note: Prompt-based polarity scores.
In table 3, I present some descriptive statistics for the different sentiment measures. Reference letter sentiment is overwhelmingly positive, regardless of how one measures it. However, there is sizeable variation in exactly how positive the referee’s opinion is. Interestingly, half of the variance is explained by within-candidate variation — different referees having differing opinions on the same candidate.
Table 3: Sentiment Polarity Descriptives
A. Descriptive Stats
Big Bird Prompt Vader LM Senticnet HGI
$N$ 1898 1898 1898 1898 1898
$mean$ 0.256 0.245 0.539 0.310 0.476
$stddev$ 0.196 0.116 0.144 0.118 0.199
$within$ 0.636 0.536 0.441 0.521 0.499
$min$0.015 -0.166 0.000 0.000 -0.089
$max$ 0.847 0.780 1.000 0.630 1.000
B. Correlations
Big Bird Prompt 1.000 0.154 0.145 0.164 0.117
Vader 1.000 0.434 0.408 0.670
LM 1.000 0.430 0.531
Senticnet 1.000 0.485
Notes: Loughran-McDonald (LM) polarity is re-scaled. “Within” is the fraction of within individual variance.
Panel B of the table presents Pearson correlations among the different sentiment measures. While all correlations are positive, it is noteworthy that the lexicon-based measures have much stronger associations among each other —correlations close to 0.5— than with the prompt-based polarity score.
Figure 3: Prompt-based Sentiment Polarity Distribution

### v2

@@ 2.3 Prompt-based Learning
Soon after the first transformers were rolled out, it became apparent to researchers that the models had capabilities that far exceeded simple language generation. The newly discovered “emergent properties” of LLMs range from text classification to translation tasks (see Brown et al., 2020). As a result of this discovery, in recent years there has been a paradigm shift away from the fine-tuning approach and toward a prompt-based learning approach. [^16] For the sake of brevity, I limit the description of the paradigm shift and of prompt-based learning to the bare minimum. See Sun et al. (2022), Min et al. (2023) and Liu et al. (2023) for comprehensive reviews. Bu et al. (2024) focus on sentiment analysis via prompt learning.
@@
Instead of fine-tuning an LLM to downstream tasks, these tasks can be reformulated to look more like those solved during the model’s pre-training. The key insight is that a textual prompt can be devised to elicit the desired prediction. For example, in order to recognize the emotion of the tweet “I missed the bus today!”, we may append the prompt “I felt so [MASK]” and ask the LLM to fill the blank with an emotion-bearing word. [^17] The same prompting strategy can be applied to multiple tasks. To elicit a translation, one might use the prompt: “English: I missed the bus today! French: …” and use a generative LLM to generate the translation.
@@
The great advantage of reformulating the problem this way is that, once the prompting strategy has been set, one may proceed to making predictions without the need of any supervised training. Specifically, prompting spares researchers from the often unsurmountable task of collecting a large labeled data set.
@@
Prompt-based learning takes advantage of the emergent properties of pre-trained model $\mathrm{P}\left(\boldsymbol{x};\boldsymbol{\theta}_{pre}\right)$ of text $\boldsymbol{x}$ itself. For sentiment analysis, the derivation of predicted probabilities $\mathrm{P}\left(y\mid\boldsymbol{x}\right)$ requires three steps:
@@
The second step of the prompting strategy involves defining a verbalizer $\mathcal{V}(\boldsymbol{m})$. Given a set of permissible answers $\mathcal{M}$, a verbalizer is just a mapping $\mathcal{V}(\boldsymbol{m})\mathrel{\mathop{\ordinarycolon}}\mathcal{M}\rightarrow\mathcal{Y}$ assigning each permissible answer to a label. In the movie review example, the permissible answers could be $\mathcal{M}=\{\text{``excellent''};\text{``good''};\text{``OK''};\text{``bad''};\text{``horrible''}\}$. And the verbalizer could be defined as follows:
$\mathcal{V}=$ $\left\{(\text{``excellent''},+);(\text{``good''},+);(\text{``OK''},\sim);(\text{``bad''},-);(\text{``horrible''},-)\right\}$
@@
The third and final step is simply to collect predicted probabilities for each label using the LLM. Let the function $f_{fill}(\boldsymbol{x}^{\prime},\boldsymbol{m})$ fill-in the location [MASK] in prompt $\boldsymbol{x}^{\prime}$ with the permissible answer $\boldsymbol{m}$. The predicted probabilities for each label are thus defined:
$\mathrm{P}(y\mid\boldsymbol{x})=$ $\sum_{\mathbf{m}\in\mathcal{M}}\mathrm{I}\left[\mathcal{V}(\boldsymbol{m})=y\right]\cdot\mathrm{P}\left(f_{fill}(f_{prompt}(\boldsymbol{x}),\boldsymbol{m});\boldsymbol{\theta}_{pre}\right),\quad y\in\mathcal{Y}$ (1)
where $\mathrm{I}[\cdot]$ is the indicator function. Intuitively, the prompt-based probabilities are a measure of how likely the LLM judges each of the answered prompts. In particular, the predicted probability of assigning positive sentiment to the input text is simply the sum of the probabilities—based on the LLM’s pre-training—assigned to each of the positive tokens in the verbalizer. This intuitive interpretation is one of the advantages of prompt-based sentiment extraction. In addition, because sentiment scores are derived from the predicted probabilities, they have a natural unit of measurement (p.p.).
@@
$\mathrm{polarity}(\boldsymbol{x})=$ $\mathrm{P}\left(y=+\mid\boldsymbol{x}\right)-\mathrm{P}\left(y=-\mid\boldsymbol{x}\right)$ (2)
In other words, prompt-based sentiment polarity is simply the difference between the sums of the probabilities that the model assigns to positive and negative completions of the prompt respectively. This intuitive interpretation greatly simplifies the use of prompt-based sentiment extraction for downstream tasks such as regression analysis.
2.3.1 RLs: Sentiment Prompt and Verbalizer
To extract sentiment from RLs, I use the template/verbalizer combination described in table 1.
Table 1: RL Prompt-based Sentiment Extraction
Prompt ($\mathcal{T}_{sentiment}$) Verbalizer ($\mathcal{V}_{sentiment}$)
“[X] In summary, this job market candidate is [MASK]” positive: {excellent, outstanding, exceptional, brilliant, great, superb, fantastic, terrific, amazing, phenomenal, awesome, stellar, wonderful, extraordinary, remarkable, impressive, incredible, fabulous, formidable, marvelous, perfect, fascinating} negative: {bad, terrible, disappointing, poor, awful, lousy, atrocious, pathetic, unsatisfactory, substandard, mediocre, common, inferior, average, ordinary, regular}
2.3.2 Model Choice
In principle, any LLM capable of evaluating $\mathrm{P}\left(\boldsymbol{x};\boldsymbol{\theta_{pre}}\right)$ can be used for prompting. However, the model choice has important implications for the overall quality and precision of the sentiment scores, as well as for other somewhat more technical aspects:
@@ 2.3.2 Model Choice
    Generative models such as those in the GPT and Llama families require a prefix prompt. In other words, the [MASK] term in the template must be the last token. [^19] Because of this restriction, the [MASK] term is generally omitted altogether when presenting the prompt. As mentioned above, the reason for this is that generative models are pre-trained using causal language modeling, i.e. predicting the next word in text sequences. In contrast, bidirectional models like BERT, distilBERT, Big Bird, or ModernBert are pre-trained via masked language modeling, so the [MASK] term can be anywhere in the template. [^20] These are called cloze prompts.
@@
    An important consideration is that the length of the input text (measured in tokens) must fit within the model’s context size, which varies from model to model. If a document is longer than the context size, it will have to be processed using some kind of rolling-window approach.
The main model I use to generate prompt-based sentiment scores is Llama 3.1, a large [^22] I use the 8 billion parameter version, which is the largest that fits my GPU. There are larger versions, which probably would yield even better results. generative language model pre-trained on a mix of publicly available online data ($\sim$15 trillion tokens). The model has a multi-lingual vocabulary encompassing 128,256 tokens. The context size—the maximum length of an input text sequence—is also 128k. For comparison purposes, in the online appendix I also present results based on prompting ModernBert, which is an encoder pre-trained on 1.7T tokens extracted from Wikipedia, the “BookCorpus”, and code databases. [^23] ModernBert is generally less capable than Llama: it has “only” 395M parameters, its vocabulary size is 50k tokens and its context size is 8,192. However, as already mentioned, the [MASK] token can be inserted anywhere when prompting a bidirectional model. In many applications, this flexibility makes creating an appropriate prompt much easier, so it is interesting to see if a bidirectional model attains reasonably comparable performance.
2.3.3 Foundational v. Instruct Models
All the LLMs mentioned so far belong to the foundational class, meaning that they have not been fine-tuned to follow human instructions or to produce output that is “aligned” with the objectives of humans. Instruct models, like Llama-instruct (Touvron et al., 2023), GPT-instruct (Ouyang et al., 2022), or DeepSeek-V3 (DeepSeek-AI et al., 2025) are “post”-trained using supervised data of question-answer pairs, as well as other techniques like reinforcement learning with human feedback, in which the model produces several alternative outputs that a human user later ranks. [^24] World famous applications like ChatGPT add a memory layer (and other features) to the otherwise stateless instruct model.
An important additional difference is that it is not possible to access the complete output—i.e. the complete probability distribution over the vocabulary—of instruct models. Instead, the model generates text output based on a sampling procedure that uses the output probabilities as weighting factors. [^25] There are additional parameters that inform the sampling process, like the ‘temperature’ parameter. While it is possible to get the probabilities associated with generated tokens from the instruct model’s output, it is not in general possible to get the probabilities of all the tokens that were not sampled. Because of this, it is not possible to calculate a polarity score such as equation (2).
Despite this loss of information, there are advantages to using instruct-type models in order to obtain the most likely label for the text input. These models allow very sophisticated prompting functions, most notably the ability to assume different “roles” and to format the model’s answers following the user’s specifications. For comparison purposes, I obtain sentiment scores from Llama 4 Instruct and DeepSeek-V3 using the following instructions:
System Prompt “You are a helpful assistant. Analyze the letter of recommendation submitted by the user and assess how positive the letter writer is regarding the ability level and future prospects of the job market candidate. Take into account that letter writers tend to be overly optimistic regarding job market candidates. Only give the maximum score to candidates deemed exceptional. Most candidates should get a medium score (around 0).”
Human Prompt “Below is a letter of recommendation for a job market candidate: [X]”
Format Instructions The only admissible output is a floating point number between -1 and 1.
It is important to reiterate that the “score” that results from these instructions is quite unlike the polarity score from equation (2). The former consists of the sequence of text tokens making up a floating point number within -1 and 1 that results from the instruct model’s sampling process. The latter is a summary measure of the whole probability distribution over the foundational model’s vocabulary.
@@ 3 Hiring Data
The analysis in this paper is based on administrative data from a leading research department in Russia during the 2013, and 2015–2021 job markets. [^26] I designate job market cycles by the year in which recruitment formally started and the job ads were posted. Unfortunately, data for the 2014 market was not archived. To protect confidentiality, I processed all RL text data myself.
As already mentioned, I only consider job market candidates who are finishing their PhD programs and have not held a full-time academic position. Overall, my data contains 1,968 letters from 645 job market candidates. In some of the analysis, I use a sample of “complete applications” consisting of candidates for whom I have at least three letters, including the letter from the main research adviser. This latter sample has 1,800 letters from 553 candidates.
The pre-processing of the reference letters’ text was kept to a minimum. I removed headers, footers, page numbers, salutations, and some obvious mistakes from faulty conversions from pdf format. I also collected personal data from the submitted CVs, as well as personal websites and LinkedIn profiles. Specifically, I have data on the candidate’s sex, nationality, prior publications and time to PhD completion. I also collect the PhD granting institution’s location and ranking. [^27] To rank Economics departments, I use Tilburg University’s ranking, which is based on publications weighted by journal impact factor and covers the whole world (https://rankings.tilburguniversity.edu/). Finally, I hand-collected the detailed job market outcome for each candidate in my data. The job market outcome corresponds to the job held by September of the following year.
In addition, I collected data on the characteristics of the letter writers: sex, affiliation, formal position, experience (years from PhD completion), and publication rank. [^28] I focus on whether the letter writer is in the top-10 or top-5 percentile of RePec’s publications ranking (https://ideas.repec.org/top/top.person.alldetail.html).
@@ 3.1 Descriptive Statistics
Mean Min Max Mean Min Max
Gender Job Market Year
Female 0.264 0 1 2013 0.118 0 1
Region of Origin 2015 0.101 0 1
Developing Country 0.093 0 1 2016 0.098 0 1
Former Soviet Union 0.209 0 1 2017 0.091 0 1
North America 0.040 0 1 2018 0.127 0 1
Europe 0.354 0 1 2019 0.172 0 1
Asia 0.305 0 1 2020 0.149 0 1
Research Field 2021 0.144 0 1
Finance 0.271 0 1 Number of RLs
Applied 0.341 0 1 1 0.028 0 1
Macroeconomics 0.236 0 1 2 0.112 0 1
Theory 0.098 0 1 3 0.642 0 1
Econometrics 0.054 0 1 4 0.219 0 1
PhD Rank PhD Region
1–25 0.251 0 1 USA 0.493 0 1
26–50 0.229 0 1 PhD Length
51–75 0.087 0 1 Seven years or more 0.096 0 1
76–100 0.073 0 1 Publications
101–150 0.144 0 1 At Least One Top Field 0.053 0 1
151–200 0.079 0 1
201+ 0.136 0 1
Notes: NN=645. The PhD ranking is from Tilburg University. I considered publications in JIE, JET, JoE, JME, JPubE, JEH, JFE, JF, Rand, JEEA, REStat, EJ, IER, all the AEJs, and the top 5.
Table 3: Descriptive Statistics — Letter Writers
@@
Title
Full Professor 1968 0.609 0 1
Associate Professor 1968 0.208 0 1
Assistant Professor 1968 0.158 0 1
Non-academic 1968 0.025 0 1
Female 1968 0.119 0 1
Experience (years since PhD granted) 1968 18.35 0 56
Ranked in Repec Top 10% 1968 0.511 0 1
Ranked in Repec Top 5% 1968 0.345 0 1
Main Research Adviser 1968 0.310 0 1
External Referee 1968 0.194 0 1
Ref. from other field 1968 0.360 0 1
Notes: There are 1,578 unique letter writers: 1,303 wrote a single letter and 275 wrote multiple letters. ‘External’ referees are not affiliated with the candidate’s university.
@@
Table 2 summarizes the main characteristics of the job market candidates in my data. There are some notable differences between my data and similar samples collected by Eberhardt et al. (2023) and Baltrunaite et al. (2024) from British and Italian institutions, respectively. Given that my institution is based in Russia, it is not surprising to see an over-representation of applicants from the former Soviet Union. [^29] Of these, 108 are Russian, 8 are Ukrainian, 5 are Armenian, 5 from the Baltic Countries, and 5 from other countries. There is also a relatively small fraction of North Americans. Just over 26% of applicants in my data are female, compared to around 30% in their samples. Another important difference is that the samples in their studies include both new PhDs and experienced academics. Because of this, the candidates in their samples have more publications.
@@
In other respects, the samples are comparable. In particular, the fraction of candidates getting their PhD from an American institution (around 50%) and the fraction in top-ranked programs are very similar.
Table 3 presents some statistics regarding the letter writers. The main observation here involves how senior most letter writers are. Over 60% of the RL were written by someone with the rank of full-professor. The average writer has over 18 years of potential experience and almost 35% of writers are in the top-5th percentile in RePec’s ranking of authors in economics. A second important point is that only about 12% of the writers are women. These statistics are roughly comparable to those found in other studies with similarly collected samples.
3.2 Reference Letter Sentiment
Figure 1 (left) shows the kernel estimate of the distribution of prompt-based polarity scores for the full sample of RLs. Several facts are apparent. Sentiment polarity is overwhelmingly positive; there are only a couple of letters with negative polarity. The distribution of sentiment is roughly symmetric, with mean sentiment of about 0.07.303030Table A.1 in the appendix has descriptive statistics for all the sentiment measures used in this paper.
There is a clear positive relationship between letter length and letter sentiment (right panel). Advisers’ letters are distinctly longer than other referees’ letters (on average, almost 33% longer). However, there is no significant “adviser effect,” i.e. there is no significant difference between the adviser’s and other writer’s sentiment independent of letter length. Frequent RL readers will not be surprised by any of these points.
Figure 1: The Distribution of Prompt-based RL Sentiment
Notes: NN=1968. Fitted line is a locally weighted regression.
Figure 2: Prompt-based Sentiment Polarity Covariates
(a) Gender
(b) Candidate Region
(c) PhD Region
(d) PhD Rank
(e) Publications
(f) Letter Writer’s Rank